OpenStudy (anonymous):

Conic sections: 4y^2+24y+16X+33=0. I'm studying on my own with library material. I need this written in standard form. I came up with: (y+3)^2=-4(x+1.5). The text has a different answer. I need to find Vertex, directrix, axis of symmetry, and focus. Because my answer is different from the text, these are off...

OpenStudy (anonymous):

Read up on completing the square

OpenStudy (anonymous):

Do you need help solving it or ?

OpenStudy (anonymous):

Thank you. I see my error now! I'd added nine to both sides rather than 36 to the right when I completed the square!

OpenStudy (anonymous):

4(y^2 + 6Y ) = -16x - 33 4(y^2 + 6Y + 9) = -16x - 33 + 36 4(y + 3)^2 = -16x + 3 (y+3)^2 = -4x + .75